# Embedded coherent structures from MHD to sub-ion scales in turbulent solar wind at 0.17 AU

### v2

@@ 7 Multiscale minimum variance analysis
Now, we consider the whole set of structures detected by integrated LIM at different time-scale ranges, see Equation (10). As we have discussed in Section 4.1, the number of structures increases toward small scales, from nearly 200 events at MHD scales, to more than $10^{4}$ events at sub-ion scales. For all these events, we study the amplitude anisotropy of the measured fluctuations via Minimum Variance Analysis (MVA, Sonnerup & Scheible, 1998). Then, we compare the observed anisotropy with the one of the model structures crossed by a spacecraft.
@@ 7.1 Observational characteristics of coherent structures
The amplitude anisotropy of the magnetic fluctuations $\delta\mathbf{B}_{j}$ of the structure along the crossing trajectory is characterised by MVA eigenvalue ratios $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{2}$. The relative amplitude $\delta B_{struct,j}/B_{0}$ is shown in color in the Figure 8. For each range of scales, the number of structures $N$ and the filling factor $P$ are shown in the legend.
Figure 8(a) gives the results of the MVA for the raw magnetic field data during 200 s time intervals around the central times $t_{0}$ of the MHD-scale coherent structures (see the discussion of the detection method in the end of Section 4.1). The MVA results for four examples analysed in detail in Section 6 are marked on the $(\lambda_{2}/\lambda_{1}$,$\lambda_{3}/\lambda_{2})$ plane with special symbols: Example-1, TD at large scales, is a black dot; Example-2, an Alfvén vortex at large scales, is a cross; Example-3, a RD at large scales, is a plus; Example-4, two neighbouring switchbacks, is a circled dot.
For a large number of events the ratio of the minimum over intermediate eigenvalues is small $\lambda_{3}/\lambda_{2}<0.5$, while intermediate over maximum variance, $\lambda_{2}/\lambda_{1}$, takes the whole range of values from 0 to 1. Among the considered models, this zone on the eigenvalue plane corresponds only to the monopole and dipole Alfvén vortices, see Figure 6. Minimum over intermediate variance, $\lambda_{3}/\lambda_{2}$, sometime takes high values ($>0.5$), as is the case for the monopole vortex, a tangential discontinuity or a magnetic hole. Values of $\lambda_{3}/\lambda_{2}$ around 0.3 and for small $\lambda_{2}/\lambda_{1}$ can be interpreted as rotational discontinuities, see Figure 6. So, the observed distribution of $\lambda_{3}/\lambda_{2}$ as a function of $\lambda_{2}/\lambda_{1}$ can be due to a superposition of different types of coherent structures. It seems that vortices are dominant, but other types of structures may also exist.
@@
Figure 8(c,d) represent the MVA results for ion and sub-ion scale structures respectively. At ion scales, the distribution is similar to what is observed in raw data, but with more cases (2028 vs 196). Sub-ion scale structures have different distribution on the MVA eigenvalue ratios. Most of the points and especially of high amplitude events are grouped closer the left side of the eigenvalue ratios plane, where $\lambda_{2}/\lambda_{1}<0.25$. But this does not exclude any of the 5 models.
Below, in Sections 7.2 and 7.3 we propose a new systematic approach to quantify the proportions of different types of structures at different scales.
@@ 7.2 Noise level estimation
We want to compare the observed distributions of $(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$ and the degree of compressibility (defined below in equation (15)) for MHD, ion and sub-ion scale structures, with the crossings of different coherent structures models, (see Section 5). The incoherent noise affects the MVA eigenvalue ratios (shown in the bottom row of the Figure 6). The greater is the ratio $\epsilon=\delta B_{noise}/\delta B_{struct}$, the closer are the $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{2}$ to 1. Therefore, we need to estimate $\epsilon$ from observations to take into account the noise in the model crossings.
For each structure at j-th scale range we calculate the ratio of the noise $\sigma_{noise,j}$ (defined in Equation (11)) to the amplitude of the structure $\delta B_{struct,j}$:
@@ 7.3 Classification
For convenience we use the notation $(r_{21},r_{32})=(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$ for MVA eigenvalue ratios. First, we investigate the presence of magnetic holes. We use two criteria to select magnetic holes: high compressibility and linear polarisation.
A coherent structure is compressible if the magnetic field modulus $|B|$ is not constant because of the parallel magnetic fluctuations of the structure. Considering the compressibility at j-th range of scales, we filter $|B|$ (as we do for fluctuations $\delta\mathbf{B}_{j}$) to define $\delta|B|$ at the scale-range $j$. The amplitude of compression associated with a coherent structure is given as $\max(|\delta|B||)_{t\in T_{struct}}$. We normalise it by $\delta B_{struct}$ to define the compressibility of the structure:
@@
First, we impose $C_{struct}>0.8$ to select strongly compressible structures and second, we delimit the zone $(r_{32}>0.6,r_{21}<0.4)$ in the MVA eigenvalue ratios plane, that is characteristic for the magnetic hole crossings, see the bottom panel of the Figure 6 (e). This zone is a bit wider than in Figure 6 (e), because for some of magnetic holes in the observational statistics, the relative noise amplitude could be higher than $\epsilon=0.1$, as used in the model. The percentage of MHD, ion and subion structures satisfying both criteria is presented in the column Magnetic hole of Table 3. We found that magnetic holes are detected only at sub-ion scales. Among sub-ion scale structures, they account for 0.4% of the cases. We will study these events in more details in a future work.
@@
Figure 9(a) show 2D histograms ($6\times 6$ bins) of distributions of the data in $(r_{32},r_{21})$–plane for observations at MHD (top), ion (middle) and sub-ion (bottom) scales. In other words, we show the probability density $P_{obs,j}$ of observations
$P_{obs,j}(r_{32},r_{21})=N_{obs,j}(r_{32},r_{21})/N_{obs,j},$ (16)
@@
where $N_{obs,j}(r_{32},r_{21})$ is the number of the observed structures in a bin, and $N_{obs,j}$ is the total number of observed structures. The index $j$ denotes the scale range.
@@
We assume that crossings of coherent structures along trajectories with different impact parameters are equally probable and we take into account the noise from the observations, with Equation (14), as explained below. Since the dipole Alfvén vortex has an angular structure, we average the results over a uniform distribution of trajectory orientations. Then, we obtain the probability density $P(r_{32},r_{21}|\text{mod})$ of MVA eigenvalue ratios for each model structure:
$P_{j}(r_{32},r_{21}|\text{mod})=N_{mod,j}(r_{32},r_{21})/N_{mod,j}$ (17)
@@
The probability distributions for 4 different models $P_{j}(r_{32},r_{21}|\text{mod})$ are shown in columns (b-e) of Figure 9. To simulate different scales, we change the level of the noise according to the estimated value at each scale, see Equation (14).
@@
The observed distribution of MVA eigenvalue ratios $P_{obs,j}$ can be expressed as the linear combination of the conditional probabilities $P_{j}(r_{32},r_{21}|\text{mod})$, determined from the models. The positive coefficients $p_{\text{mod}}$ reflect the probability to encounter each model structure. Coefficients $p_{\text{mod}}$ are found from the constrained minimization problem:
$\begin{cases}||P_{\text{obs}}(r_{32},r_{21})-\displaystyle\sum_{\text{mod}}p_{\text{mod}}\,P(r_{32},r_{21}|\text{mod})||\to 0\\ \displaystyle\sum_{\text{mod}}p_{\text{mod}}\leq 1\\ p_{\text{mod}}\geq 0\end{cases}$ (18)
Figure 9: Probability distributions on the MVA eigenvalue ratios plane $(r_{32},r_{21})=(\lambda_{3}/\lambda_{2},\lambda_{2}/\lambda_{1})$. The column (a) shows the probability [%] per bin to observe a coherent structure with the corresponding MVA eigenvalue ratios (so $P_{obs,j}(r_{32},r_{21})$ defined by Equation (16)). First and second panels of the column (a) show the distributions for the MHD-scale coherent structures using the raw (non-filtered) data and the MHD-range filtered data respectively. Third and fourth rows of the column (a) correspond to coherent structures detected at ion and sub-ion scale ranges respectively. Columns (b-e) show the probability densities obtained from simulating model crossings ($P_{j}(r_{32},r_{21}|\text{model})$ defined by Equation (17)). The difference between panels of different rows (b-e) is due to the different imposed noise level $\epsilon_{sim}$ (see Section 7.2 for details).
@@
We use the least squares minimization. The problem is solved for each range of scales. For convenience the index $j$ is omitted in Equation 18.
@@
The resulting probabilities $p_{\text{mod}}$ are shown in the Table 3. For each range of scales the observations are most consistent with the crossings of the dipole Alfvén vortices ($\geq 80\%$). The monopole vortices account for $7-15\%$ of coherent structures among different scales. The rotational discontinuities are observed in raw (non-filtered) data at MHD scales only, but not in the MHD band-pass filtered data. We explain it by the fact that the band pass filter makes the waveforms very different from the current sheet simple model, see Figure 7(b). Tangential discontinuities does not appear to be statistically significant. There is $6\%$ of events which were not possible to model at MHD scales and $39\%$ at sub-ion scales (see the None column in Table 3). These unidentified large number of events at sub-ion scales is probably due to a more 3-dimentional nature of the fluctuations not taken into account by nearly incompressible models with mostly $\delta B_{\perp}$ fluctuations.
@@
The result presented in Table 3 does not change qualitatively if instead of least squares, the sum of the absolute values of probability differences (between observations and models) in each bin is minimized.
@@
Consider the obtained results in comparison with the previous observations. The visual classification of ion-scale coherent structures at 0.17 au, during the first PSP perihelion, has been done recently in Perrone et al. (2020). Three different time intervals were considered: quiet, weekly-disturbed and highly-disturbed solar wind. The highly-disturbed interval (of 1.5 h) with $B_{R}$ reversals is a subset of the 5h-interval considered here. The authors concluded that in the highly-disturbed interval current sheets were dominant ($46$%), while during the weekly-disturbed interval Alfvén vortices ($45$%) and wave packets ($50$%) were observed. This is in contrast with the obtained quantitative classification results at ion scales (showing that Alfvén vortices are dominant).
@@
In the previous studies of ion scales coherent structures at 1 au in slow (Perrone et al., 2016) and fast (Perrone et al., 2017) solar wind with Cluster satellites, the dominance of Alfvén vortices with respect to current sheets has been found. That is more consistent with our results at 0.17 au in the slow wind.
@@
We have visually analysed 196 MHD-scale coherent structures to understand why the percentage of current sheets is low. We remind that, the central times of the structures here are determined by the integrated LIM over MHD range between 1 and 100 s. The time interval for each structure is ±$\pm 100$ s around the central time, the corresponding spatial scale is $7\cdot 10^{4}$ km. The results are provided in the Table 4: We realise that the observed events need to be separated in isolated and non-isolated structures, like a train of Alfvén vortices (see the second column ‘Multiple Alfven vortices’) or non-isolated current sheets (column 5) and as well we observe a co-existance of a large scale weak current sheet with a vortex inside (column 3). So, in comparison with an automatic classification (Table 3), here we observe more current sheets, most of them are associated with a vortex (51 cases out of 196), but as well 6 isolated current sheets which have been detected in Raw data but not at MHD scales. For the non-isolated current sheets and especially those assosiated with a vortex, the automatic classification prosess tends to interpret the structures as Alfvén vortices. Indeed, if the current sheet is not isolated, the perturbations of the neighboring structure affect the MVA eigenvalue ratios, so that the structure is shifted away from the characteristic zone on the $(r_{21},r_{32})$ plane, see Figure 6. Consequently, these events contribute to $p_{monopole}$ and $p_{dipole}$ in the solution of the minimisation problem (Eq. 18).
@@
Finally, for the 196 events, we have visually checked as well ion and sub-ion scale sub-structures. Their shape is consistent with vortex crossings and not with current sheets. That is in agreements with Table 3.
@@ 8 Conclusion and discussions
In this paper, for the first time, we apply a multi-scale approach in physical space, from the largest MHD scales $\sim 10^{5}$ km, to the smallest resolved sub-ion scales $\sim 3$ km.
@@
Using PSP merged magnetic field data at 0.17 au and the Morlet wavelet transform, we detect intermittent coherent structures covering all scales. We find localised magnetic fluctuations on subion scales with amplitudes up to $\delta B/B_{0}\sim 0.03$, where $B_{0}$ is the local magnetic field strength. This small scale structure is typically embedded in a larger structure at ion scales, with amplitudes up to $\delta B/B_{0}\sim 0.3$. At its turn, ion scale structure is embedded in a high-amplitude ($\delta B_{struct}/B_{0}\sim 0.4-1.0$) MHD-scale structures. Such embedding across the whole turbulent cascade is presented here for the first time.
@@
The topology and properties of the coherent structures change from scale to scale. Using plasma and field time profiles, we could characterize several events in more details. We show examples of planar tangential (and rotational) discontinuities at MHD scales with the thickness of about $100\rho_{i}$ containing embedded sub-structures inside it: incompressible ion-scale Alfvén vortex (with a cross-section of $\simeq 30\rho_{i}$) and sub-ion scale compressible vortex $\simeq 5\rho_{i}$, see the sketch in Figure 10(a). Another example is an Alfvén vortex at MHD scales with the cross-section of $5\cdot 10^{3}\rho_{i}$, with embedded incompressible vortices of $60\rho_{i}$ at ion and of $\sim 10\rho_{i}$ at sub-ion scales, see the sketch in Figure 10(b).
@@
We completed study of examples with a statistical analysis. In a time interval of about 5 hours we detected nearly 200 events at the MHD scales, and much more events at ion ($\sim 2\cdot 10^{3}$) and sub-ion scales ($\sim 10^{4}$). The filling factor of the structures, that we estimate in a conservative way [^1] Using only the time where the integrated LIM is over the threshould means that the lower-energy part of each event is not taken in-to account in the calculation of the filling factor. If the method of (Perrone et al., 2016) was used, the filling factor would be 2-3 times larger. (see discussion in section 4.1), decreases from 12% at MHD scales to 7% and 6% at ion and sub-ion scales correspondingly.
@@
To determine the dominant type of coherent structures, we perform an automatic classification based on the comparison of the observed amplitude anisotropy of magnetic fluctuations within the observed events at all scales with analytical models of Alfvén vortices, current sheets and magnetic holes. We do not consider magnetosonic shocks as far as the analysed time interval is mostly incompressible at MHD scales. The results show the dominance of Alfvén vortices at all scales and only few current sheets mostly in raw data.
@@
In order to understand the low number of current sheets, we did a visual inspection of 196 events at MHD scales. It reveals that isolated current sheets are indeed rare (3%). Most of the events are rather complex. About 10% of structures represent non-isolated current sheets. Many of the detected events are vortices within the current sheets (26%). Isolated and non-isolated vortices account for 66% of the structures. In all these cases, except isolated current sheets, the automatic classification method tend to interpret the data as vortices.
@@
Ion transition range of scales corresponds to $[30,340]$ km, or $[5,70]\rho_{i}$ in our time interval. Here, with the automatic classification method we found 85% dipole vortices and 15% of monopoles. Planar discontinuities are not found by our method. Visual inspection of 196 cases confirms the absence of current sheets at these scales.
@@
On sub-ion scales ($[3,30]$ km or $[0.5,5]\rho_{i}$) coherent structures represent dipole vortices (49%), monopole vortices (7%), rotational current sheets (5%) and magnetic holes (0.4%). Around 39% of sub-ion scale structures do not fit any of the considered models. It is plausibly, because the incompressible models of vortices have been used in comparison to observations. To improve this study at these scales in the future, the compressible ion scales Alfvén vortex model of (Jovanović et al., 2020) should be used. In this model, together with typical vortical $\delta B_{\perp}$, the compressible component $\delta B_{\|}$ appears, which is in pressure balance with density fluctuations. We presume, that the comparison with this model will increase the proportion of the vortices at sub-ion scales.
@@
Results presented in this article show the dominance of Alfvén vortices at all scales. Thus, Alfvén vortices are important building blocks of solar wind turbulence from inertial range to sub-ion scales. This may explain $\delta B/B_{0}=\xi\delta V/V_{A}$ with $\xi\neq 1$ typically observed in the solar wind for Alfvénic periods.
@@
Our results are limited to a specific slow highly-perturbed solar wind region at 0.17 au from the Sun. The analysis can be expanded to different solar wind conditions (different radial distance, types of solar wind, originating from ecliptic or polar regions of the Sun) in order to obtain a more general picture.
@@
Multiscale nature of coherent strucutres described in this article can be studied in future by the Helioswarm (future NASA mission). It will cover MHD, ion and sub-ion scales at the same time.

### v4

@@ 7 Multiscale minimum variance analysis
Now, we consider the whole set of structures detected by integrated LIM at different time-scale ranges, see Equation (11). As we have discussed in Section 4.1, the number of structures increases toward small scales, from nearly 200 events at MHD scales to more than $10^{4}$ events at sub-ion scales. For all these events, we study the amplitude anisotropy of the measured fluctuations via Minimum Variance Analysis (MVA, Sonnerup & Scheible, 1998). Then, we compare the observed anisotropy with one of the model structures crossed by a spacecraft.
@@ 7.1 Observational characteristics of coherent structures
The amplitude anisotropy of the magnetic fluctuations $\delta\mathbf{B}_{j}$ of the structure along the crossing trajectory is characterized by MVA eigenvalue ratios $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{2}$. The relative amplitude $\delta B_{struct,j}/B_{0}$ is shown in color in the Figure 8. For each range of scales, the number of structures $N$ and the filling factor $P$ (defined in the same way as eq. (10) but for each scale range) are shown in the legend.
Figure 8(a) gives the results of the MVA for the raw magnetic field data during 200 s time intervals around the central times $t_{0}$ of the MHD-scale coherent structures (see the discussion of the detection method at the end of Section 4.1). The MVA results for four examples analyzed in detail in Section 6 are marked on the $(\lambda_{2}/\lambda_{1}$,$\lambda_{3}/\lambda_{2})$ plane with special symbols: Example-1, TD at large scales, is a black dot; Example-2, an Alfvén vortex at large scales, is a cross; Example-3, a RD at large scales, is a plus; Example-4, two neighbouring switchbacks, is a circled dot.
For a large number of events, the ratio of the minimum over intermediate eigenvalues is small $\lambda_{3}/\lambda_{2}<0.5$, while intermediate over maximum variance, $\lambda_{2}/\lambda_{1}$, takes the whole range of values from 0 to 1. Among the considered models, this zone on the eigenvalue plane corresponds only to the monopole and dipole Alfvén vortices, see Figure 6. Minimum over intermediate variance, $\lambda_{3}/\lambda_{2}$, sometimes takes high values ($>0.5$), as is the case for the monopole vortex, a tangential discontinuity, or a magnetic hole. Values of $\lambda_{3}/\lambda_{2}$ around 0.3 and for small $\lambda_{2}/\lambda_{1}$ can be interpreted as rotational discontinuities, see Figure 6. So, the observed distribution of $\lambda_{3}/\lambda_{2}$ as a function of $\lambda_{2}/\lambda_{1}$ can be due to a superposition of different types of coherent structures. It seems that vortices are dominant, but other types of structures may also exist.
@@
Figure 8(c,d) represent the MVA results for ion and sub-ion scale structures, respectively. At ion scales, the distribution is similar to what is observed in raw data, but with more cases (2028 vs 196). Sub-ion scale structures have different distributions on the MVA eigenvalue ratios. Most of the points (especially yellow ones, corresponding to high-amplitude events) are grouped closer to the left part of the eigenvalue plane, where $\lambda_{2}/\lambda_{1}<0.25$. But this does not exclude any of the 5 models.
Below, in Sections 7.2 and 7.3, we propose a new systematic approach to quantify the proportions of different types of structures at different scales.
@@ 7.2 Noise level estimation
We want to compare the observed distributions of $(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$ and the degree of compressibility (defined below in equation (15)) for MHD, ion and sub-ion scale structures, with the crossings of different coherent structures models, (see Section 5). The incoherent noise affects the MVA eigenvalue ratios (shown in the bottom row of Figure 6). The greater is the ratio $\epsilon=\delta B_{noise}/\delta B_{struct}$, the closer are the $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{2}$ to 1. Therefore, we need to estimate $\epsilon$ from observations to take into account the noise in the model crossings.
For each structure at j-th scale range, we calculate the ratio of the noise $\sigma_{noise,j}$, defined in Equation (12), to the amplitude of the structure $\delta B_{struct,j}$:
@@ 7.3 Classification
For convenience we use the notation $(r_{21},r_{32})=(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$ for MVA eigenvalue ratios. First, we investigate the presence of magnetic holes. We use two criteria to select magnetic holes: high compressibility and linear polarization.
A coherent structure is compressible if the magnetic field modulus $|\mathbf{B}|$ is not constant because of the parallel magnetic fluctuations of the structure. Considering the compressibility at j-th range of scales, we filter $|\mathbf{B}|$ (as we do for fluctuations $\delta\mathbf{B}_{j}$) to define $\delta|\mathbf{B}|$ at the scale-range $j$. The amplitude of compression associated with a coherent structure is given as $\max(|\delta|\mathbf{B}||)_{t\in T_{struct}}$. We normalise it by $\delta B_{struct}$ to define the compressibility of the structure:
@@
First, we impose $C_{struct}>0.8$ to select strongly compressible structures and second, we delimit the zone $(r_{32}>0.6,r_{21}<0.4)$ in the MVA eigenvalue ratios plane, that is characteristic for the magnetic hole crossings, see the bottom panel of the Figure 6 (e). This zone is a bit wider than in Figure 6 (e), because for some of the magnetic holes in the observational statistics, the relative noise amplitude could be higher than $\epsilon=0.1$, as used in the model. The percentage of MHD, ion, and subion structures satisfying both criteria is presented in the column Magnetic hole of Table 3. We found that magnetic holes are detected only at sub-ion scales. Among sub-ion scale structures, they account for 0.4% of the cases. We will study these events in more detail in future work.
@@
Figure 9(a) shows 2D histograms ($6\times 6$ bins) of distributions of the data in $(r_{21},r_{32})$–plane for observations at MHD (top), ion (middle) and sub-ion (bottom) scales. In other words, we show the probability density $P_{obs,j}$ of observations
$P_{obs,j}(r_{21},r_{32})=N_{obs,j}(r_{21},r_{32})/N_{obs,j},$ (17)
@@
where $N_{obs,j}(r_{21},r_{32})$ is the number of the observed structures in a bin, and $N_{obs,j}$ is the total number of observed structures. The index $j$ denotes the scale range.
@@
We assume that crossings of coherent structures along trajectories with different impact parameters are equally probable and we take into account the noise from the observations, with Equation (15), as explained below. Since the dipole Alfvén vortex has an angular structure, we average the results over a uniform distribution of trajectory orientations. Then, we obtain the probability density $P(r_{21},r_{32}|\text{mod})$ of MVA eigenvalue ratios for each model structure:
$P_{j}(r_{21},r_{32}|\text{mod})=N_{mod,j}(r_{21},r_{32})/N_{mod,j}$ (18)
@@
The probability distributions for 4 different models $P_{j}(r_{21},r_{32}|\text{mod})$ are shown in columns (b-e) of Figure 9. To simulate different scales, we change the noise level according to the estimated value at each scale, see Equation (15).
@@
The observed distribution of MVA eigenvalue ratios $P_{obs,j}$ can be expressed as the linear combination of the conditional probabilities $P_{j}(r_{21},r_{32}|\text{mod})$, determined from the models. The positive coefficients $p_{\text{mod}}$ reflect the probability to encounter each model structure. Coefficients $p_{\text{mod}}$ are found from the constrained minimization problem:
$\begin{cases}||P_{\text{obs}}(r_{21},r_{32})-\displaystyle\sum_{\text{mod}}p_{\text{mod}}\,P(r_{21},r_{32}|\text{mod})||\to 0\\ \displaystyle\sum_{\text{mod}}p_{\text{mod}}\leq 1\\ p_{\text{mod}}\geq 0\end{cases}$ (19)
@@
where ||.||||.||| | . | | means a norm (the square root of the square of the difference of the matrix) that allow us to use the least squares minimization. The problem is solved for each range of scales. For convenience, the index $j$ is omitted in Equation (19).
Figure 9: Probability distributions on the MVA eigenvalue ratios plane $(r_{21},r_{32})=(\lambda_{2}/\lambda_{1},\lambda_{3}/\lambda_{2})$. The column (a) shows the probability [%] per bin to observe a coherent structure with the corresponding MVA eigenvalue ratios (so $P_{obs,j}(r_{21},r_{32})$ defined by Equation (17)). The first and second panels of column (a) show the distributions for the MHD-scale coherent structures using the raw (non-filtered) data and the MHD-range filtered data respectively. The third and fourth rows of column (a) correspond to coherent structures detected at ion and sub-ion scale ranges. Columns (b-e) show the probability densities obtained from simulating model crossings ($P_{j}(r_{32},r_{21}|\text{model})$ defined by Equation (18)). The difference between panels of different rows (b-e) is due to the different imposed noise levels $\epsilon_{sim}$ (see Section 7.2 for details).
@@
The resulting probabilities $p_{\text{mod}}$ are shown in the Table 3. For each range of scales, the observations are most consistent with the crossings of the dipole Alfvén vortices ($\geq 80\%$). The monopole vortices account for a small fraction of coherent structures, $7-15\%$ depending on the scale range. The rotational discontinuities are observed in raw (non-filtered) data at MHD scales only, but not in the MHD band-pass filtered data. This is because the band-pass filter makes the waveforms very different from the current sheet simple model, see Figure 7(b). Tangential discontinuities do not appear to be statistically significant. There is $6\%$ of events which were not possible to model at MHD scales and $39\%$ at sub-ion scales (see the None column in Table 3). These unidentified large number of events at sub-ion scales is probably due to a more 3-dimensional nature of the fluctuations not taken into account by nearly incompressible models with mostly $\delta B_{\perp}$ fluctuations.
@@
The result presented in Table 3 does not change qualitatively if, instead of least squares, the sum of the absolute values of probability differences (between observations and models) in each bin is minimized.
@@
Let us compare the obtained results with the previous observations. The visual classification of ion-scale coherent structures at 0.17 au, during the first PSP perihelion, has been done recently in Perrone et al. (2020). Three different time intervals were considered: quiet, weekly-disturbed, and highly-disturbed solar wind. The highly-disturbed interval (of 1.5 h) with $B_{R}$ reversals is a subset of the 5h-interval considered here. The authors concluded that in the highly-disturbed interval current sheets were dominant ($46$%), while during the weekly-disturbed interval Alfvén vortices ($45$%) and wave packets ($50$%) were observed. This is in contrast with the quantitative classification results obtained here at ion scales, showing that Alfvén vortices are dominant.
@@
In the previous studies of ion scales coherent structures at 1 au in slow (Perrone et al., 2016) and fast (Perrone et al., 2017) solar wind with Cluster satellites, the dominance of Alfvén vortices with respect to current sheets has been found. This is more consistent with our results at 0.17 au in the slow wind.
@@
We have visually analyzed 196 MHD-scale coherent structures to understand why the percentage of current sheets is low. We remind that the peaks of the integrated LIM over the MHD range (between 1 and 100 s) determine the central times of the structures. The time interval for each structure is ±$\pm 100$ s around the central time, the corresponding spatial scale is $\sim 7\cdot 10^{4}$ km. The results are summarized in the Table 4: We realized that the observed events need to be separated in isolated and non-isolated structures, like a train of Alfvén vortices (see the second column ‘Multiple Alfven vortices’) or non-isolated current sheets (column 5) and as well we observe a co-existence of a large scale weak current sheet with a vortex inside (column 3). So, in comparison with an automatic classification (Table 3), here we observe more current sheets, most of them are associated with a vortex (51 cases out of 196), but as well 6 isolated current sheets which have been detected in Raw data but not at MHD scales. For the non-isolated current sheets and especially those associated with a vortex, the automatic classification tends to interpret the structures as Alfvén vortices.
Indeed, if the current sheet is not isolated, the perturbations of the neighboring structure affect the MVA eigenvalue ratios, so that the structure is shifted away from the characteristic zone on the $(r_{21},r_{32})$–plane, see Figure 6. Consequently, these events contribute to $p_{monopole}$ and $p_{dipole}$ in the solution of the minimisation problem (Eq. 19).
@@
Finally, for the 196 events, we have visually inspected ion and sub-ion scale sub-structures. Their shape is consistent with vortex crossings and not with current sheets. This is in agreement with Table 3.
@@ 8 Conclusion and discussions
In this paper, for the first time, we apply a multi-scale approach in physical space, from the largest MHD scales $\sim 10^{5}$ km to the smallest resolved sub-ion scales $\sim 3$ km.
@@
Using PSP merged magnetic field data at 0.17 au and the Morlet wavelet transform, we detect intermittent coherent structures covering all scales. We find localized magnetic fluctuations on subion scales with amplitudes up to $\delta B/B_{0}\sim 0.03$, where $B_{0}$ is the local magnetic field strength. These small-scale structures are typically embedded in a larger structure at ion scales, with amplitudes up to $\delta B/B_{0}\sim 0.3$. At its turn, the ion scale structure is embedded in a high-amplitude ($\delta B_{struct}/B_{0}\sim 0.4-1.0$) MHD-scale structure. Such embedding across the whole turbulent cascade is presented here for the first time.
@@
The topology and properties of the coherent structures change from scale to scale. Using plasma and magnetic field time profiles, we characterize several events in more detail. We show examples of planar tangential and rotational discontinuities at MHD scales with the thickness of about $100\rho_{i}$ containing embedded sub-structures inside it: incompressible ion-scale Alfvén vortex (with a cross-section of $\simeq 30\rho_{i}$) and sub-ion scale compressible vortex $\simeq 5\rho_{i}$, see the sketch in Figure 10(a). Another example is an Alfvén vortex at MHD scales with the cross-section of $5\cdot 10^{3}\rho_{i}$, with embedded incompressible vortices of $60\rho_{i}$ at ion and of $\sim 10\rho_{i}$ at sub-ion scales, see the sketch in Figure 10(b).
@@
We completed the study of examples with a statistical analysis. In a time interval of about 5 hours we detected nearly 200 events at the MHD scales, and many more events at ion ($\sim 2\cdot 10^{3}$) and sub-ion scales ($\sim 10^{4}$). The filling factor of the structures, that we estimate in a conservative way [^1] Using only the time where the integrated LIM is over the threshold means that the lower-energy part of each event is not taken in-to account in the calculation of the filling factor. If the method of (Perrone et al., 2016) was used, the filling factor would be 2-3 times larger. (see discussion in section 4.1), decreases from 12% at MHD scales to 7% and 6% at ion and sub-ion scales correspondingly.
@@
To determine the dominant type of coherent structures, we perform an automatic classification based on the comparison of the observed amplitude anisotropy of magnetic fluctuations within the observed events at all scales with analytical models of Alfvén vortices, current sheets, and magnetic holes. We do not consider magnetosonic shocks as far as the analyzed time interval is mostly incompressible at MHD scales. The results show the dominance of Alfvén vortices at all scales and only a few current sheets, and mostly in raw data.
@@
In order to understand the low number of current sheets, we did a visual inspection of 196 events at MHD scales. It reveals that isolated current sheets are indeed rare (3%). Most of the events are rather complex. About 10% of structures represent non-isolated current sheets. Many of the detected events are vortices within the current sheets (26%). Isolated and non-isolated vortices account for 66% of the structures. In all these cases, except isolated current sheets, the automatic classification method tends to interpret the data as vortices.
@@
The ion transition range of scales corresponds to $[30,340]$ km, or $[5,70]\rho_{i}$ in our time interval. Here, with the automatic classification method, we found 85% dipole vortices and 15% of monopoles. Planar discontinuities are not found by our method. Visual inspection of 196 cases confirms the absence of current sheets at these scales.
@@
On sub-ion scales ($[3,30]$ km or $[0.5,5]~{}\rho_{i}$) coherent structures represent dipole vortices (49%), monopole vortices (7%), rotational current sheets (5%) and magnetic holes (0.4%). Around 39% of sub-ion scale structures do not fit any of the considered models. It is, possibly, because the incompressible models of vortices have been used to compare with the observations. To improve this study at these scales in the future, the compressible ion scales Alfvén vortex model of (Jovanović et al., 2020) should be used. In this model, together with typical vortical $\delta\mathbf{B}_{\perp}$, the compressible component $\delta B_{\|}$ appears, which is in pressure balance with density fluctuations. We presume, that the comparison with this model will increase the proportion of the vortices at sub-ion scales.
@@
Results presented in this article show the dominance of Alfvén vortices at all scales : from inertial to sub-ion range. Thus, Alfvén vortices are important building blocks of solar wind turbulence and in particular of its intermittency at all scales. Alfvén vortices can explain $\delta\mathbf{V}_{\perp}/V_{A}=\xi\,\delta\mathbf{B}_{\perp}/B_{0}$, with $\xi\neq 1$, typically observed in the solar wind for Alfvénic periods.
In the reflection-driven turbulence in the reduced magnetohydrodynamic numerical simulation of Meyrand et al. (2023), the authors observe cellularisation of turbulence with generation of magnetic vortices with $\delta\mathbf{V}_{\perp}=0$, with $k_{\|}=0$ and with magnetic field discontinuity at vortex boundary. These structures are thus quite different from the smooth Alfvén vortices observed here with $\delta\mathbf{B}_{\perp}\sim\delta\mathbf{V}_{\perp}$ correlations, but they have similar twisted magnetic field configuration and $k_{\|}=0$.
@@
Our results are limited to a specific slow highly-perturbed solar wind region at 0.17 au from the Sun. The analysis can be expanded to different solar wind conditions (different radial distances, types of solar wind, originating from ecliptic or polar regions of the Sun) to obtain a more general picture.
@@
The multiscale nature of coherent structures described in this article can be studied in the future by the Helioswarm (NASA mission). It will cover MHD, ion, and sub-ion scales at the same time.